OpenStudy (anonymous):

Simplify (7x2– 2x + 9) + (3x2+ 15x – 4) A. 10x2+ 13x + 5 b. 10x2– 13x – 5 c. 10x2+ 17x + 13 d. 10x2– 17x – 13

hartnn (hartnn):

combine like terms like 7x^2+3x^2=(7+3)x^2=10x^2 can u do the same for other terms??

OpenStudy (anonymous):

-2x + 15x = 13x

OpenStudy (anonymous):

9 - 4 = 5

OpenStudy (anonymous):

@hartnn would it be A?

hartnn (hartnn):

yup :)